1.) A circle of area 24 square units has a sector of area 10 square units. .the central angle formed by the sector
1
Expert's answer
2013-01-09T08:52:46-0500
Area of a circle / area of a sector=24/10=2.4
Then forcentral angles we have the same 2*pi/x=2.4 then x=2*pi/2.4=pi/1.2 radian=5pi/6
Hence central angle is 5pi/6